4 and 5 in first pic
C in sec pic
C in third pic
e in fourth pic

Part I A Table 3.3 Calculation of x2 for a sample removed from a large population consisting of equal number of colored and white beans Classes Observed Expected Deviations OE) (O-EVE (phenotypes) (O) (O-E) Colored 0.6 White | Totals 15 1. How many degrees of freedom do you have in the interpretation of the x2 value? 2. Using table 3.2, determine what x values lie on either side of the xi calculated in Table 3.3. record values here: 1o4 27. b 3. What probability values are associated with those table values of x?? 4. In the space provided, write a brief interpretation of the x value you have just calculated 5. Place the x2 value for your sample in the proper column (under the appropriate P value) of the table on the chalkboard that is prepared in the style of Table 3.4.
media%2F6af%2F6af8e0fe-4faa-4355-baa6-b7
media%2F0d8%2F0d81cc8c-0eaa-44a1-af8e-ab
media%2Fecc%2Feccaec6b-0d4a-4e4e-a4cb-19
Answer #1

Answer 4:

The difference between the observed and the expected values for each category is 1.2

Answer 5:

?^2 = 1.2

Answer c:

Degrees of freedom, df = n - 1 = 6 - 1 = 5

Using chi-square tables, the probability is

P [X^2 (5) > 8.0843] = 0.1517

Answer c:

Degrees of freedom, df = n - 1 = 4 - 1 = 3

Using chi-square tables, the probability is

P [X^2 (3) > 3.4732] = 0.3243

Answer e:

Degrees of freedom, df = 2 - 1 = 1

Using chi-square tables, the p-value is

P [X^2 (1) > 656.51] = 0.000

Reject Ho